Triangle YPQ is isosceles with base PQ. If YQ=5y+6, QP=12y+1.2, and YP=10y-11, find the length of each of each side of triangle YPQ.
Answer by stanbon(75887)

Triangle YPQ is isosceles with base PQ. If YQ=5y+6, QP=12y+1.2, and YP=10y-11, find the length of each of each side of triangle YPQ.
---------
If PQ is the base the other two sides must be equal::
Equation:
5y+6 = 10y-11
-----
5y = 17
y = 17/5
----
YQ = 5y+6 = 5(17/5)+6 = 23
QP = 12y+1.2 = 12(17/5)+1.2 = 40.8+1.2 = 50
YP = 10y-11 = 10(17/5)-11 = 23
=====================
Cheers,
Stan H.
=====================